64 identical drops of mercury are charged simultaneously to the same potential of 10 volt. assuming that the drops to be spherical, if all the charged drops are made to combine to form one large drop, then its potential will be

Grade:12th pass

1 Answers

Vikas TU
7 years ago
64*(4/3 pi r^3) = (4/3 pi R^3)
R = 4r
Potential of bigger sphere,
k64q/R => 16(kq/r) = >16*10 = 160 volt.